OpenStudy (anonymous):

For the position function s(t) =-16t^2 + 90t + 60, make a table of average velocities and then make a conjecture about the instantaneous velocity at t=4

OpenStudy (anonymous):

Time Interval [3,4] [3.9,4] [4,5] [4,4.1]

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what's the avg velocity from t = 3 to t = 4?

jimthompson5910 (jim_thompson5910):

any ideas?

OpenStudy (anonymous):

nope

jimthompson5910 (jim_thompson5910):

remember the formula m = (f(x2) - f(x1))/(x2 -x1) we used earlier?

OpenStudy (anonymous):

so do i 3-4/3-4?

jimthompson5910 (jim_thompson5910):

f(x) = -16x^2 + 90x + 60 so f(3) = ???

OpenStudy (anonymous):

186?

jimthompson5910 (jim_thompson5910):

now what is f(4)

OpenStudy (anonymous):

164

OpenStudy (anonymous):

is the answer 22?

jimthompson5910 (jim_thompson5910):

more like -22

OpenStudy (anonymous):

ok i'm getting it

jimthompson5910 (jim_thompson5910):

AROC = (f(x2) - f(x1))/(x2 - x1) AROC = (f(4) - f(3))/(4 - 3) AROC = (164 - 186)/(4 - 3) AROC = (-22)/(1) AROC = -22 Note: AROC stands for Average Rate Of Change

OpenStudy (anonymous):

Jim...your the best!!!

jimthompson5910 (jim_thompson5910):

glad to be of help